OpenStudy (anonymous):

The factors for x2 + 7x – 44 are x - 4 and x + 11 ??

OpenStudy (lgbasallote):

let's try... (x-4)(x+11) x^2 + 11x - 4x - 44 x^2 + 7x - 44 you got it!

OpenStudy (anonymous):

Maybe you should take a break, @0mgl0la. You've been at this for several hours now.

OpenStudy (anonymous):

WHYY??

OpenStudy (lgbasallote):

too much studying isnt good for you lol

OpenStudy (anonymous):

LOL.

OpenStudy (anonymous):

\[x ^{2}+7x-44=0\] \[x ^{2}+11x-4x-44=0\] \[x(x+11)-4(x+11)=0\] \[(x+11)(x-4)=0\]